If you want to estimate the proportion of all RIT students who smoke within 0.05 with 90% confidence, what is the minimum sample size you will need?

a. 1562

b. 2033

c. 271

d. 549
Answer #1

Answer: c. 271

Confidence interval for P = ( ľ F Ê CHÊ Margin of Error = P( 1 ) * 2012 { s use 8=0.5. it maximizes Ø 11-p) x = 0.10 n 0.05 =